OpenStudy (anonymous):

how to determine whether y varies directly with x? and if it does how would i find the constant of variation k and write the equation

OpenStudy (anonymous):

equations that are similar to the linear function y=mx + b. y = kx k = y/x. K is the slope of the linear function i think.

OpenStudy (anonymous):

but how do i figure out if y varies directly with x

OpenStudy (anonymous):

well, if u have an equation that is similar to the linear function, y = mx + b, then you know that the shape of that equation is linear meaning that y has a direct relationship with x. In other words, you know that y varies with x.

OpenStudy (anonymous):

for example, newton's second law , Fnet = ma. As you can see, its y varies directly with x.